# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ A.3 Ablation experiments on the ASR module
A.3.1 ASR models evaluation with oracle diarization
@@
Table 10 shows the tcpWERs of our developed ASR backends when using Oracle diarization and our SE frontend. The tcpWERs of all systems were significantly improved compared to those of the baseline. Although the best single ASR system is the Whisper Large v3 (ASR1), the WavLM transducer (ASR4), which is the smallest system, achieved comparable performance. Note that we trained two versions of ASR1 and ASR4 with different training settings, ASR1’ and ASR4’ in Table 10, respectively. For ASR1, two models were built with and without curriculum learning, as detailed in Section 5.2.1. As seen in Table 10, we see that our proposed curriculum learning improved recognition performance on all datasets. For ASR4, two models were trained with and without SpecAugment [86] during the fine-tuning step using only the CHiME-8 dataset. The two models exhibited close tcpWERs, slightly better when using SpecAugment. ASR5 consists of the combination of 12 hypotheses, which are 1-best hypotheses generated by the six ASR systems with and without LM restoring.
@@
We observe that the four ASR models exhibit different error trends. ASR1 achieves a lower deletion error rate compared to all other systems. This may be due to the curriculum learning scheme, which filters out unreliable utterances that may include errors in GSS preprocessing, where the speaker’s speech is removed but the label remains assigned, resulting in the model improving the recognition of more words. For insertion and substitution errors, ASR3 and ASR4 perform better in terms of insertions, while ASR3 and ASR1 do better in terms of substitutions.
@@ A.3.2 Impact of LM rescoring on performance
To evaluate the impact of LM rescoring, we modified the hypotheses used in ROVER and conducted a performance comparison. As we discussed in Section 5.3, we applied LM rescoring to the recognition results of each ASR model and then performed system combination using ROVER with the original ASR hypotheses and the rescored hypotheses.

### v2

@@ A.3 Ablation experiments on the ASR module
A.3.1 evaluation of ASR models with Oracle diarization
@@
Table 10 shows the tcpWERs of our developed ASR backends when using Oracle diarization and our SE frontend. The tcpWERs of all systems were significantly improved compared to those of the baseline. Although the best single ASR system is the Whisper Large v3 (ASR1), the WavLM transducer (ASR4), which is the smallest system, achieved a comparable performance. Note that we trained two versions of ASR1 and ASR4 with different training settings, ASR1’ and ASR4’ in Table 10, respectively. For ASR1, two models were built with and without curriculum learning, as detailed in Section 5.2.1. As seen in Table 10, our proposed curriculum learning improved the recognition performance on all datasets. For ASR4, two models were trained with and without SpecAugment [90] during the fine-tuning step using only the CHiME-8 dataset. The two models exhibited close tcpWERs, slightly better when using SpecAugment. ASR5 consists of the combination of 12 hypotheses, which are 1-best hypotheses generated by the six ASR systems with and without LM restoring.
@@
We observe that the four ASR models exhibit different error trends. ASR1 achieves a lower deletion error rate compared to all other systems. This may be due to the curriculum learning scheme, which filters out unreliable utterances that may include errors in GSS preprocessing, where the speaker’s speech is removed but the label remains assigned, resulting in the model improving the recognition of more words. As for insertion and substitution errors, ASR3 and ASR4 perform better in terms of insertions, while ASR3 and ASR1 do better in terms of substitutions.
@@ A.3.2 Impact of LM rescoring on performance
To evaluate the impact of LM rescoring, we modified the hypotheses used in ROVER and conducted a performance comparison. As discussed in Section 5.3, we applied LM rescoring to the recognition results of each ASR model and then performed system combination using ROVER with the original ASR hypotheses and the rescored hypotheses.